UT 20160320 LUCI1 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, APramskiy, WSeifert, DThompson,
Observer AO: DMiller, EPinna (remote from Tucson if required), APuglisi (remote from Italy, GTaylor
Telescope Operator: DHuerta
LBTO Support: TSargent

Plan for the night

- Switch to LUCI1-AO Com

- Do various tests from ComPlan with Bins=1-4, most importantly stability of system

Summary

- ND-filters for Halogen lamps installed into LUCI2 CalUnit, countrates are OK now

- More progress by Doug & Greg on the FLAO performance

- LUCI1- AO Com has really started now. Several scripts were succesfully executed. During a given script the loop opened in about 66% of the cases

- IQ down to 0.08" on coadded images (20min) with Bin = 2 achieved

Details

All times are UT
Files are luci1.20160320.NNNN.fits

Walter & Elliot put in 0.7ND-filters for the 3 halogen lamps in the CalUnit of LUCI2 to dimm light a bit
=> Both halogen lamps deliver now a similar amount of light (see below)


from Dave's logs: ---

These should be LUCI2 CU data!

N3.75 imaging
FieldStop not recentered, @N30 position
LIR:NORM 1x2.75s each

Halo3 ON
0001 J 4160
0002 H 6350
0003 K 21700
0004 Ks 10900
0005 z 640

what is 0006?

## Halo2 ON Halo3 OFF
0007 z 15400 ADU in 2.75s
0008 HeI 3050
0009 Pgam 2450
0010 Pbeta 5000
0011 FeII 12300
0012 H2 12250
0013 BrG 13300
0014 OH1060 1900
0015 OH1190 3300
0016 Y1 8000
0017 Y2 13150
0018 Jlo 34100 - too bright
- was actually 0021 ... check headers for correct filters & lamps

## Halo2 OFF Halo3 ON
0022 Jhi 1580
0023 Jlo 1420
FW1: clear, blind and PVlens not done

## G200 N1.8 LS_1.00arcsec
0024 zJspec@1.17 1x2.75s Halo3 51 ADU
0025 --"-- 1x20s 250 ADU
0026 --"-- 1x2.75s Halo2 750
0027 --"-- 1x30s 8000
0028 --"-- 1x2.75s Halo1 3800
0029 HKspec@1.93 1x2.75s Halo1 17000
0030 --"-- 1x10s Halo2 13450

### G210 N1.8 LS_1.00arcsec
0031 K @ 2.20um 1x10s Halo2 3100
0032 H @ 1.65um 1x10s Halo2 2450
0033 J @ 1.25um 1x10s Halo2 1000
0034 J @ 1.25um 1x10s Halo1 4950
0035 z @ 1.00um 1x10s Halo1 1900

### N30 Imaging, N30_FieldStop in FPU
0036 - forgot to change camera!
0037 z 1x10s Halo3 ~25 ADU/pixel
0038 --"-- Halo2 ~570
0039 --"-- Halo1 2870 ???
0042 J 1x2.75s Halo1 5120
0043 H 1x2.75s Halo1 7570
0044 Ks 1x2.75s Halo1 6000
0045 K 1x2.75s Halo1 8520
0046 K 1x2.75s Halo2 2040
0047 H --"-- 1650
0048 Ks 1390
0049 J 1000
0050 dark 1x2.75s
0051 dark 1x10s
0052 dark 1x10s again

#########################################
### LUCI1 N30 imaging
luci1.20160320.NNNN.fits

0001 K all lamps off ~180 ADU
0002 K 1x2.75s Halo3 290 ADU/pixel
0003 --"-- Halo2 2925
0004 H 1x2.75s Halo2 2890
0005 J 1x2.75s Halo2 1830
0006 Ks 1x2.75s Halo2 2080
0007 z 1x2.75s Halo2 440
0008 --"-- Halo1 11800
0009 1x2.75s dark


End of DTompson's logs ---

01:35 Sunset Sky is clear (very few cirrus around), T = +4 deg, Hum = 25%, low wind

Dave is checking and centering Pupil


From Daves logs ---

0010 0019 images taken during adjustment
Adjustments:
FM1m1 -4750 Y FM1m2 +5300 X
FM4m1 -5500 Y FM4m2 +5400 X

Switch to N30 camera to position N30_FieldStop
FM4m1 4600, -540
FM4m2 9900, -9900, -9900 ... close enough


End of Dave's logs ---

01:52 David checks pointing and collimates

Stability of system, accuracy of offsets

02:14 Preset to p16_0220_20mar2016 (Psyche (16)), OK, seeing 0.9 on DImm, DThompson aborted collimation #0024, RunAO failed,
resent preset, didn't work, restarted scheduler, no star on AO, restart GCS, preset p16_0240_20mar2016, OK; loop closed,
bin 2, 153 modes, #0025-0032, images saturated in K-band, resent preset, this time with Br_gam filter, p16BrG_0300_20mar2016, OK,
loop closed, #0033-0035, first ime w/o NCPA, stopped because of absolute offsets in script (always referring to the FIRST position on the detector)
Typo in script (guide star pmdec said "gstrPmDec")

Stability of PSF

03:15 Preset to Oj287_a, failed, David checks pointing, resent preset, OK, object only 15deg from moon,lets see how it goes, OK
started with bin=4 since RS-mag in script was from catalogue and off by 2mag (RS is Blazar who varies by definition)
moved automaticlaly to bin = 3, 66 modes, no NCPA, loop closed, DIMM=1.2, cirrus passing, IQ 0.12",aberration on 1st ima, #0037-0041

03:33 Preset to OJ287_PSFa, bin = 3, 66 modes, #0042-0046, ima qual 0.12"

03:46 Preset to Oj287_b, changed mag of RS-star (Blazar) in script to the one determined by Greg, and yes, started with bin = 3,
ima qual 0.11", #0047-0051

03:58 Preset to OJ287_PSFb, bin = 3, 66 modes, #0052-56, star close to egde in x, need to check coords

Check all images, all are elongated, possibly vibrations, mots likely imperfect optical gain

Ima quality as fct. of pos on detector

04:12 Preset to M82, preset failed, David checked pointing, resent preset, OK, bin = 2, 153 modes, cirrus in, 1mag absorption,
gosh, RS-star has 9.5mag, loop closed, Ks-band: #0057-0058, loop opened just before end of exposure, loop closed again, loop opened
Did preset again, loop closed, loop opened again, KS-band, ##0060- , cirrus getting thicker, loop opened again, star has faded by 2.5mag,

04:40 ALADIN hangs and than all OBSGUI, xkill, restarted, OK, reason was old Aladin version on obs2, new version installed

05:05 Optical gain can not follow that quick for that change of extinction, stopped

BIN=1 tests

Doug and Greg taking over, doing some tests with BIN=1 and high elevation since seeing is good
Other tests do not make sense due to high extinction

AFC-tests

06:50 Greg and Doug are done with their tests. Still heavy Cirrus and now degrading seeing (about 1.4"),
Alexander & Dave start with some AFC-tests

Using K band, 10s interval, SRR 200s, Nread=21

Spots should be a couple hundred ADU above the bckgd

@ EL=87, meridian crossing...AFC does not keep image completely stable.

@ EL=65, 1.5h east...AFC still showing some image-to-image shift

Files: LUCI1 0099 0103, LUCI2 0066 0070

@ EL~65 Turn off FlexComp but keep telescope/rotator tracking

Files: LUCI1 0104 0108, LUCI2: 0071 0075

* Worse residuals than with AFC, as expected

Still @ EL~65 turn off telescope & rotator tracking

Files: LUCI1 0109 0111, LUCI2 0076 0078

* No net flexure between the last two images, as expected

08:45 Cirrus start to disappear, seeing imporves, continue...

System performance, PSF-stability

08:48 Preset to 3C294, Bin=2, 153 modes, loop closed, RS-star is double with 0.15" separation!!!
Dimm reports 0.7", gnagna, #0116-0120, mistake in script, remade script and restartet,
preset OK, runAO failed (stage timeout), #0121-0127, loop opened, modified script again, new PA,
target coincides with crosstalk features by the RS on previous images, Ok, #00128-0150,
stopped, cirrus came back

Bin=1-tests, cont'd

Doug and Greg taking over agian, doing some tests with BIN=1 and high elevation since seeing is good
Images look MUCH better now !!!

Test for dynamic range

11:51 Preset to NGC6543, OK, RunAO, Bin=2, 153 modes, Ks, Br_gam, #0181+, beautiful images, Dimm reports 0.8"
loop opens at #0204 (skip frame), cont'd, some more Br_gam, last image #209

12:50 AO-stopped, too bright (30min before sunrise)

13:00 David is closing up

Lessons learned

  • Scriptor reliably delivers scripts for AO-observing

  • Avoid to put your science target north/south of the RS in DETXY, otherwise crosstalk from the RS will kill your science

  • Difraction spikes are much less of a concern as eg on Keck

-- %USERSIG{JochenHeidt - 2016-03-20}%
Topic revision: r10 - 09 Jun 2016, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback